OpenStudy (anonymous):

Geometry help please! Prove the Triangle Exterior Theorem. The measure of each exterior angle of a triangle equals the sum of the measurements of its two remote interior angles. Question in comments below.

OpenStudy (anonymous):



OpenStudy (cwrw238):



OpenStudy (cwrw238):

m <1 + m < 4 = 180 degrees (angles on same line) m<2 + m<3 + m<4 = 180 ( 3 angles of triangle add up to 180 degrees) since m < 1 plus m<4 = 180 and m<2 + m<3 plus m< 4 is also = 180 degrees m<1 = m <2 + m < 3

OpenStudy (anonymous):

Awesome! Thank you for explaining it :)

OpenStudy (cwrw238):

yw - that is how i was taught in UK - the US method might use different terms.